OpenStudy (anonymous):

g(x) = 5x2 - 4x h(x) = 3x + 9 Find: g(h(x))

OpenStudy (anonymous):

\[45^2+15x+45\]

OpenStudy (anonymous):

which can be simplified into: \[15(3x^x+x+3)\]

OpenStudy (hoblos):

g(h(x)) = 5(3x+9)^2 -4(3x+9) = 5( 9x^2 + 54x +81) - 12x -36 = 45x^2 + 270x + 405 -12x -36 = 45x^2 + 234x + 269

OpenStudy (anonymous):

^ he's right...

OpenStudy (anonymous):

I'm being really stupid today... '-.-

OpenStudy (hoblos):

sorry it 369 instead of 269

OpenStudy (anonymous):

thanks both of you

OpenStudy (hoblos):

45x^2 + 234x + 369

OpenStudy (hoblos):

my pleasure XD